# Asymptotically Unbiased Synthetic Control Methods by Density Matching

### v4

@@ 10 Conclusion
We have highlighted the problem of implicit endogeneity in the linear regression models commonly employed by existing SCMs. To address this issue, we proposed assuming mixture models among outcome distributions and developed a density-matching SCM. By matching densities, our method consistently estimates SC weights and reinforces the theoretical foundation for using SCMs in comparative studies. We demonstrated the effectiveness of this approach through both simulation experiments and empirical applications.
@@
First, we confirm the basic facts about the uniqueness of the CHF.
Proposition A.1 (Theorem 26.2 in Billingsley (1995)).
Distinct measures cannot have the same characteristic function.
Let us define
$\widehat{Q}_{G}(\bm{w})\coloneqq\sum_{\gamma\in\{1,2,\dots,G\}}\frac{2h^{\gamma+1}}{(\gamma+1)!}\left|\frac{1}{T_{0}}\sum_{t\in\mathcal{T}_{0}}\left(\Big{(}Y^{N}_{j,t}\Big{)}^{\gamma}-\sum_{j\in\mathcal{J}^{U}}w_{j}\Big{(}Y^{N}_{j,t}\Big{)}^{\gamma}\right)\right|,$
$Q_{G}(\bm{w})\coloneqq\sum_{\gamma\in\{1,2,\dots,G\}}\frac{2h^{\gamma+1}}{(\gamma+1)!}\left|\mathbb{E}_{0,t}\left[\Big{(}Y^{N}_{j,t}\Big{)}^{\gamma}\right]-\sum_{j\in\mathcal{J}^{U}}w_{j}\mathbb{E}_{j,t}\left[\Big{(}Y^{N}_{j,t}\Big{)}^{\gamma}\right]\right|.$
@@
First, from the definition of $\Delta^{J}$, $\Delta^{J}$ is a compact parameter space. Next, $\widehat{Q}(\bm{w})$ converges uniformly to $Q(\bm{w})$ as $T_{0}\to\infty$ under Assumption 6.1. Furthermore, $Q(\bm{w})$ is continuous with respect to $\bm{w}$.
@@
    $\widehat{Q}_{G}\left(\widehat{\bm{w}}^{\mathrm{DMSCM}}\right)>\widehat{Q}(\bm{w}_{0})-\varepsilon/3$;
@@
    $\widehat{Q}_{G}(\widehat{\bm{w}}^{\mathrm{DMSCM}})>\widehat{Q}_{G}\left(\widehat{\bm{w}}^{\mathrm{DMSCM}}\right)-\varepsilon/3$;
@@
    $\widehat{Q}_{G}\left(\bm{w}_{0}\right)>Q_{G}(\bm{w}_{0})-\varepsilon/3$.
@@
$Q_{G}\left(\widehat{\bm{w}}^{\mathrm{DMSCM}}\right)>\widehat{Q}_{G}\left(\widehat{\bm{w}}^{\mathrm{DMSCM}}\right)-\varepsilon/3>\widehat{Q}_{G}\left(\bm{w}_{0}\right)-2\varepsilon/3>Q_{G}\left(\bm{w}_{0}\right)-\varepsilon.$
Thus, for any $\varepsilon>0$, $Q_{G}\left(\widehat{\bm{w}}^{\mathrm{DMSCM}}\right)>Q_{G}\left(\bm{w}_{0}\right)-\varepsilon$ with probability approaching one. Let $\mathcal{N}$ be any open subset of $\Delta^{J}$ containing $\widetilde{\Phi}^{*}$. Since $\Theta\cap\mathcal{N}^{c}$ is compact, and $Q_{G}(\bm{w})$ is continuous with respect to $\bm{w}$, we have
$\sup_{\bm{w}\in\Delta^{J}\cap\mathcal{N}^{c}}Q_{G}(\bm{w})=Q_{G}(\bm{w}^{\dagger})<Q_{G}(\bm{w}_{0})$
for some $\bm{w}^{\dagger}\in\Theta\cap\mathcal{N}^{c}$. Thus, choosing $\varepsilon=Q_{G}(\bm{w}_{0})-\sup_{\bm{w}\in\Delta^{J}\cap\mathcal{N}^{c}}Q_{G}(\bm{w})$, it follows that
$Q_{G}\left(\widehat{\bm{w}}^{\mathrm{DMSCM}}\right)>\sup_{\bm{w}\in\Delta^{J}\cap\mathcal{N}^{c}}Q_{G}(\bm{w}),$
and hence $\widehat{\bm{w}}^{\mathrm{DMSCM}}\in\mathcal{N}$, where recall that $\mathcal{N}$ is any open subset of $\Delta^{J}$ containing $\widetilde{\Phi}^{*}$. Thus, $\widehat{\bm{w}}^{\mathrm{DMSCM}}_{T_{0},G}\xrightarrow{\mathrm{p}}\bm{w}^{*}_{G}\in\widetilde{\Phi}^{\dagger}$ holds as $T_{0}\to\infty$. Additionally, from Proposition A.1, we have $\bm{w}^{*}_{G}\to\bm{w}^{*}\in\widetilde{\Phi}^{*}$ as $G\to\infty$. ∎

### v5

@@ 8 Conclusion
This study investigated the asymptotically unbiased estimation of the treatment effect in the setting of SCMs. We first pointed out the problem of implicit endogeneity in the linear regression models commonly employed by existing SCMs, which can be interpreted as a form of measurement error bias. To address this issue, we focused on distributional SCMs, in which we assume mixture models for the outcome distributions and developed a moment-matching SCM. By matching the moments between the outcome of the treated unit and the weighted sum of the outcomes of the untreated units, our method consistently estimates SC weights and reinforces the theoretical foundation for using SCMs in comparative studies. We demonstrated the effectiveness of this approach through both simulation experiments and empirical applications.
@@
This section provides the proof for Theorem 5.8. Before presenting the proof, we introduce some existing results related to the moment problem and additional notations.
Moment problem. The question of whether a sequence of moments uniquely determines a distribution is known as the moment problem (Schmüdgen, 2020). In particular, when the random variable is bounded, the problem is referred to as the Hausdorff moment problem. For this case, Hausdorff (1921) provides the following result: if a random variable has support on $[0,1]$, then the sequence of its moments is uniquely determined by its distribution.
Proposition A.1 (Hausdorff (1921), Satz III).
Let $\{\mu_{n}\}_{n\geq 0}$ be a real sequence. Define the forward differences recursively by
$\Delta\mu_{n}\coloneqq\mu_{n+1}-\mu_{n},\qquad\Delta^{k}\mu_{n}\coloneqq\Delta(\Delta^{k-1}\mu_{n})\quad(k\geq 1).$
The following are equivalent:
  • (Moment representation) There exists a finite positive Borel measure $\nu$ supported on $[0,1]$ such that
    $\mu_{n}=\int^{1}_{0}u^{n}d\nu(u)\qquad(n=0,1,2,\dots).$
  • (Complete monotonicity) The sequence is completely monotone, that is
    $(-1)^{k}\Delta^{k}\mu_{n}\geq 0\qquad\text{for all }k,n\in\mathbb{N}.$
Moreover, when these conditions hold, the representing measure $\nu$ is unique.
Additional notations. Let us define
$\widehat{Q}_{G}(\bm{w})\coloneqq\sum_{\gamma\in\{1,2,\dots,G\}}v_{\gamma}\left|\frac{1}{T_{0}}\sum_{t\in\mathcal{T}_{0}}\left(\Big(Y^{N}_{j,t}\Big)^{\gamma}-\sum_{j\in\mathcal{J}^{U}}w_{j}\Big(Y^{N}_{j,t}\Big)^{\gamma}\right)\right|,$
$Q_{G}(\bm{w})\coloneqq\sum_{\gamma\in\{1,2,\dots,G\}}v_{\gamma}\left|\mathbb{E}_{0,t}\left[\Big(Y^{N}_{j,t}\Big)^{\gamma}\right]-\sum_{j\in\mathcal{J}^{U}}w_{j}\mathbb{E}_{j,t}\left[\Big(Y^{N}_{j,t}\Big)^{\gamma}\right]\right|.$
Proof. Then, we prove Theorem 5.8 as follows:
@@
First, from the definition of $\Delta^{J}$, $\Delta^{J}$ is a compact parameter space. Next, $\widehat{Q}(\bm{w})$ converges uniformly in probability to $Q(\bm{w})$ as $T_{0}\to\infty$ under Assumptions 5.4 and 5.7 (Pötscher & Prucha, 1989). Furthermore, $Q(\bm{w})$ is continuous with respect to $\bm{w}$.
@@
    $\widehat{Q}_{G}\left(\widehat{\bm{w}}^{\mathrm{MMSCM}}\right)<\widehat{Q}(\bm{w}_{0})+\varepsilon/3$;
@@
    $Q_{G}(\widehat{\bm{w}}^{\mathrm{MMSCM}})<\widehat{Q}_{G}\left(\widehat{\bm{w}}^{\mathrm{MMSCM}}\right)+\varepsilon/3$;
@@
    $\widehat{Q}_{G}\left(\bm{w}_{0}\right)<Q_{G}(\bm{w}_{0})+\varepsilon/3$.
@@
$Q_{G}\left(\widehat{\bm{w}}^{\mathrm{MMSCM}}\right)<\widehat{Q}_{G}\left(\widehat{\bm{w}}^{\mathrm{MMSCM}}\right)+\varepsilon/3<\widehat{Q}_{G}\left(\bm{w}_{0}\right)+2\varepsilon/3<Q_{G}\left(\bm{w}_{0}\right)+\varepsilon.$
Thus, for any $\varepsilon>0$, $Q_{G}\left(\widehat{\bm{w}}^{\mathrm{MMSCM}}\right)<Q_{G}\left(\bm{w}_{0}\right)+\varepsilon$ holds with probability approaching one. Let $\mathcal{N}$ be any open subset of $\Delta^{J}$ containing $\widetilde{\Phi}^{*}$. Since $\Delta^{J}\cap\mathcal{N}^{c}$ is compact, and $Q_{G}(\bm{w})$ is continuous with respect to $\bm{w}$, we have
$\inf_{\bm{w}\in\Delta^{J}\cap\mathcal{N}^{c}}Q_{G}(\bm{w})=Q_{G}(\bm{w}^{\dagger})>Q_{G}(\bm{w}_{0})$
for some $\bm{w}^{\dagger}\in\Delta^{J}\cap\mathcal{N}^{c}$. Thus, choosing $\varepsilon=Q_{G}(\bm{w}_{0})-\inf_{\bm{w}\in\Delta^{J}\cap\mathcal{N}^{c}}Q_{G}(\bm{w})$, it follows that
$Q_{G}\left(\widehat{\bm{w}}^{\mathrm{MMSCM}}\right)<\inf_{\bm{w}\in\Delta^{J}\cap\mathcal{N}^{c}}Q_{G}(\bm{w}),$
with probability approaching one. This implies that $\widehat{\bm{w}}^{\mathrm{MMSCM}}\notin\Delta^{J}\cap\mathcal{N}^{c}$, and hence $\widehat{\bm{w}}^{\mathrm{MMSCM}}\in\mathcal{N}$, where recall that $\mathcal{N}$ is any open subset of $\Delta^{J}$ containing $\widetilde{\Phi}^{*}$. Therefore, $\widehat{\bm{w}}^{\mathrm{MMSCM}}_{T_{0},G}\xrightarrow{\mathrm{p}}\bm{w}^{*}_{G}\in\widetilde{\Phi}^{\dagger}$ holds as $T_{0}\to\infty$. Additionally, from Proposition A.1, we have $\bm{w}^{*}_{G}\to\bm{w}^{*}\in\widetilde{\Phi}^{*}$ as $G\to\infty$. ∎
Appendix B Proof of Theorem 5.12
Proof.
Each draw $Y^{l,\dagger}_{0}$ is independent and distributed according to $\widehat{F}_{Y^{N}_{0,t},T_{0}}(y)$. Therefore, for any fixed $T_{0}$, we have
$G_{T_{0},L}\xrightarrow{\mathrm{a.s.}}\widehat{F}_{Y^{N}_{0,t},T_{0}}(y)$
as $B\to\infty$, from the Glivenko–Cantelli theorem.
Next, from Theorem 5.8, $\widehat{\bm{w}}^{\mathrm{MMSCM}}_{T_{0},G}\xrightarrow{\mathrm{p}}\bm{w}^{*}_{G}\in\widetilde{\Phi}^{\dagger}$ holds as $T_{0}\to\infty$.
Lastly, $F_{Y^{N}_{j,t},T_{0}}(y)$ converges uniformly in probability to $F_{Y^{N}_{0,t}}(y)$ as $T_{0}\to\infty$ under Assumptions 5.4 and 5.7 (Pötscher & Prucha, 1989).
By combining above results, we have
$\sup_{y}\big|G_{T_{0},L}(y)-F_{Y^{N}_{0,t}}(y)\big|\xrightarrow{\mathrm{p}}0\quad(L\to\infty,n\to\infty),$
Appendix C Confidence bands
In Tables 24 and Figures 79, we present the results of the conformal inference analysis described in Section 7.2. We omitted MMSCM ($G=100$) due to the space limitation.